# Ferraz, Cappart, and Vidal

### v1

@@
Deep Learning for Data-Driven Districting and Routing
@@
Department of Computer and Software Engineering, Polytechnique Montréal, Canada
@@
SCALE-AI Chair in Data-Driven Supply Chains
Department of Mathematics and Industrial Engineering, Polytechnique Montréal, Canada
Department of Computer Science, Pontifical Catholic University of Rio de Janeiro
@@
Districting-and-Routing, Strategic Optimization, Supervised Learning, Graph Neural Networks
@@ 1 Introduction
Districting is the process of partitioning a service region, represented as a collection of basic geographical units, into larger clusters called districts. This practice is ubiquitous in large-scale transportation and last-mile delivery systems for mail delivery (Bruno et al. 2021), home care services (Benzarti, Sahin, and Dallery 2013), and maintenance services (García-Ayala et al. 2016). A delivery policy in fixed districts has several benefits: (1) allowing the separation and the aggregation of the requests in advance before all information is available, (2) reducing the complexity of the task thanks to the decomposition of the routing optimization process, (3) stimulating the familiarity of drivers and thus their efficiency within their respective geographical regions (Zhong, Hall, and Dessouky 2007), and (4) increasing the satisfaction of customers thanks to a higher familiarity with their drivers (Kovacs et al. 2014).
@@
Districting decisions are strategic and linked with major financial and societal stakes. These decisions typically hold for months or years, whereas operational routes occur on a daily or weekly basis and are subject to variations. Optimizing or even evaluating districting decisions is a very complex task. Demands are uncertain and volatile, and routing cost evaluations typically translate into large-scale vehicle routing problems, which are time-consuming to solve and highly sensitive to the spatial distribution of the requests. Because of these two different classes of decisions and planning horizons, the related districting-and-routing problems still represent important challenges (Drexl and Schneider 2015, Kalcsics and Ríos-Mercado 2019).
@@
A common method to estimate costs on a long planning horizon is to consider each district (e.g., a few dozen), use a Sample Average Approximation approach (SAA – Verweij et al. 2003) to generate sample demand scenarios (e.g., 10 to 500) within it, and solve a routing problem for each scenario and district. This approach offers the benefit of obtaining a good estimation, but at the cost of extensive computational time due to the large number of districts and scenarios. It is typically applicable to a fixed districting solution but becomes impractical within search algorithms for districting problems that require evaluating numerous districting solutions (e.g., a local search). Therefore, most approaches for districting-and-routing rely on continuous approximation formula to evaluate costs (Franceschetti, Jabali, and Laporte 2017). It is, for instance, the case of the Beardwood-Halton-Hammersley (BHH) formula for the traveling salesman problem (Beardwood, Halton, and Hammersley 1959), which is frequently applied to estimate routing costs through $n$ independently distributed points in a compact area of size $A$ as $\alpha\sqrt{nA}$, where $\alpha$ is a constant. Though extremely fast, this evaluation approach is far less accurate than a scenario-based approach and may even drive the search toward solutions that have sub-optimal characteristics (see Section 6.4).
@@
Against this background, we leverage the recent progress in deep learning algorithms, specifically Graph Neural Networks (GNN – Scarselli et al. 2009, Kipf and Welling 2017), to accurately model expected routing costs and investigate their impact within solution methods for strategic districting-and-routing. Predictions with this approach should be fast enough to be used for solution-cost evaluations within a local search algorithm, and accurate enough to drive the search toward near-optimal districting solutions. Geographical areas are divided into geographical units called Basic Units (BU) and represented as a network of BUs with edges linking contiguous areas. Therefore, a solution to the districting problem corresponds to a partition of this graph into connected components representing contiguous districts. Our GNN learns to approximate the routing costs associated with any connected component assimilated to a district. It is trained on a set of districts with known costs obtained using SAA, i.e., by solving TSPs for demand scenarios with Lin–Kernighan algorithm (Lin and Kernighan 1973) and taking their average value as the expected cost.
@@
The trained GNN gives us an oracle that estimates routing costs within districts accurately, as demonstrated by numerical experiments using out-of-sample districts. Furthermore, we must validate if this performance translates into better strategic districting decisions, since even precise cost estimators may perform badly when integrated within an optimization process due to skew or outliers. To obtain a complete picture, we analyze the impact of using the proposed GNN within search algorithms for districting-and-routing. To that end, we design a prototypical solution approach for districting, which creates initial districts using a mathematical programming method for graph partitioning, and then subsequently applies an Iterated Local Search (ILS – Lourenço, Martin, and Stützle 2019) to generate better districting plans. The solution approach can use the GNN or any other baseline approach as a cost oracle, so we can conduct extensive computational experiments to compare the impact of different cost estimators on the structure and quality of the final solutions. Overall, the contributions of this paper are fourfold:
@@
    We design an ILS that leverages the predicted routing costs to generate high-quality districting plans within minutes. The construction of the initial solution is done by solving an adapted flow formulation for the balanced connected $k$-partition problem, whereas the local search and perturbation procedures are based on BU relocations and exchanges. Our method is generic and can use any cost oracle; it can produce efficient districting plans for practitioners and allows us to gauge the performance of accurate routing-cost estimates in the solution process.
@@
    We conduct extensive numerical experiments using the data of five metropolitan areas in the United Kingdom (Bristol, Manchester, Leeds, London, and West-Midlands), which contain up to 120 basic geographical units. We compare different cost oracles regarding the quality of routing cost estimates and their performance on the districting task, with districts of different sizes, instances of different scales, and different depot configurations. As seen in these experiments, the proposed GNN-based approach achieves a much smaller prediction error on average compared to other baselines, allowing long-term economic gains of $10.12\%$ on average (and exceeding $20\%$ in some cases) due to better districting decisions.
@@
    We investigate why cost estimation accuracy has a much larger impact than initially anticipated on the districting-and-routing problem. Indeed, whereas cost estimation accuracy differs by one or two percentage points between methods, their impact on districting-and-routing solution costs is one order of magnitude larger. A finer-grained analysis of the districting solutions shows that classical cost estimation methods tend to drive the search toward suboptimal solution structures and that the compactness of the districts is not sufficient to warrant good districting-and-routing solutions in all situations.
@@ 2 Related Works
Districting is a broad class of problems arising from diverse applications related to politics, service, sales, and distribution, among others (Kalcsics and Ríos-Mercado 2019). Political districting is especially important in democracies where each district elects a parliament member. Horn, Hampton, and Vandenberg (1993) compares different compactness measurements in the context of politics, while Webster (2013) presents a review of different districting measurements and their socio-economical impacts. Another review of Zoltners and Sinha (2005) is focused on districting for sales territory design. The variety of applications has also led to different solution methodologies. Bozkaya, Erkut, and Laporte (2003) introduced a tabu search metaheuristic to solve a multi-criteria districting problem considering compactness and population equality, among other objectives. Benzarti, Sahin, and Dallery (2013) introduced and solved different MILP formulations for home health care (HHC) districting, while García-Ayala et al. (2016) proposed an arc-based approach for problems with an underlying network structure, including postal delivery, meter readings, winter gritting, road maintenance, and municipal solid waste collection. Bruno et al. (2021) solved the problem of reorganizing postal delivery services in Bologna. Lei et al. (2015) studied a multiple traveling salesmen and districting problem with multiple periods and depots for sales territory applications. They introduced an adaptive large neighborhood search considering multiple objectives related to dissimilarity, equity, and profit. Finally, for distribution districting, Novaes, de Cursi, and Graciolli (2000) solved a vehicle fleet and districting optimization problem using an initial ring-radial topology and optimized it with a genetic algorithm. Galvão et al. (2006) extended this work using Voronoi diagrams to refine districts. Zhong, Hall, and Dessouky (2007) provided a tabu search heuristic that considers drivers’ familiarity to estimate routing costs. Finally, Carlsson (2012) used geometric arguments to draw district boundaries and obtain a balanced workload.
Long-term routing cost estimations in the context of districting. The objectives considered in districting depend on the application context and may include compactness, workload balance, service levels, and costs, among others. Our study considers a districting-and-routing problem in which the districts are designed to optimize delivery costs on an extended planning horizon. Estimating precisely those costs with SAA requires solving routing problems for a large number of scenarios. Although relatively practical when evaluating a single districting solution, this approach becomes a considerable bottleneck when many solutions must be evaluated and compared. To address this issue, several works have replaced SAA-based cost estimations with surrogate measures that require less intensive computation. A widespread way to estimate routing costs (Daganzo 1984, Rosenfield, Engelstein, and Feigenbaum 1992, Novaes and Graciolli 1999, Novaes, de Cursi, and Graciolli 2000, Galvão et al. 2006, Lei, Laporte, and Guo 2012, Lei et al. 2015) consists of a parametrized weighted sum of a line-haul factor (e.g., smallest or average distance from the depot to the district) and an intra-district routing distance as evaluated by BHH (Beardwood, Halton, and Hammersley 1959). Some other works (Benzarti, Sahin, and Dallery 2013, Bruno et al. 2021) have also taken compactness into account when estimating the travel distance.
Beardwood-Halton-Hammersley theorem states that the optimal routing cost of $R$ randomly distributed points within an area $A$ is equal to $\alpha\sqrt{RA}$ where $\alpha$ is a positive constant. This formula has been further extended by Daganzo (1984) for vehicle routing problems, leading to the following estimation: $2L+0.57\sqrt{RA}$, where $L$ is the average distance from the depot to any point inside the district. Along this line, Chien (1992) performed Monte Carlo simulations to determine the best constant for different estimators. Perhaps the first use of machine learning for approximating routing costs has been proposed by Kwon, Golden, and Wasil (1995). They construct regression models and shallow neural networks to estimate the optimal length of a traveling salesman tour of customers located in a rectangular region. Because of computational limitations, this work focused on instances with up to 80 customers, using a shallow neural network containing one hidden layer with three neurons. Akkerman and Mes (2022) developed different learning models to estimate routing costs based on known customer positions, and tested the methods on a rolling horizon to support customer selection decisions. Finally, other continuous approximation formulas have been presented in the broader literature on districting (Figliozzi 2007, Franceschetti, Jabali, and Laporte 2017) but, to our best knowledge, no study has used any other formula than BHH to design distribution districts.
Machine learning approaches for routing cost estimations. The last decade has witnessed the rise of deep learning architectures (LeCun, Bengio, and Hinton 2015). Thanks to the increasing computational power and available data, deeper and more sophisticated neural network architecture achieved exceptional performances on various tasks, from computer vision (Krizhevsky, Sutskever, and Hinton 2012) to natural language processing (Vaswani et al. 2017) and games (Silver et al. 2017, Wang et al. 2021). On the lookout for new applications, the machine learning community has progressed on the design of deep learning-based methods for combinatorial optimization problems (Mirhoseini et al. 2021, Bengio, Lodi, and Prouvost 2021). For instance, Bello et al. (2016) introduced a reinforcement learning approach to solve the traveling salesman problem, and was able to solve approximately problems up to a hundred nodes. Deudon et al. (2018) relied on an attention mechanism (Bahdanau, Cho, and Bengio 2015, Vaswani et al. 2017) to improve the solving process and used a 2-OPT mechanism to quickly improve the results. Attention has also been used by Kool, van Hoof, and Welling (2018), who proposed an extended model able to solve different routing problems. Joshi, Laurent, and Bresson (2019) introduced a GNN with beam search to solve the traveling salesman problem, achieving an average gap of $1.39\%$ for 100 nodes, improving the learning mechanisms to solve the problem but still remaining far from standard optimization approaches. Still, Joshi et al. (2022) have shown that despite the progress of learning-based approaches, cases with more than 100 nodes can still not be successfully solved, and the authors argue that the lack of generalization of current models may require rethinking neural combinatorial optimization pipeline.
More recently, GNNs have gained popularity for solving approximately routing problems. Similar to convolutional neural networks, which are dedicated to learning from spatial data such as images, GNNs are specialized to learn from data having a graph structure, as in many combinatorial problems related to routing (Xie et al. 2019, Rusek et al. 2019, Derrow-Pinion et al. 2021, Kafaei et al. 2021, 2023). There are nowadays many architectures based on graph neural networks, e.g., with attention (Veličković et al. 2018), with gates (Li et al. 2016), or with sampling (Hamilton, Ying, and Leskovec 2017). An extensive survey of graph neural networks and their use in combinatorial optimization can be found in Cappart et al. (2023).
Interestingly, despite the trend pioneered by Kwon, Golden, and Wasil (1995) and the considerable progress of deep learning architectures, no sophisticated machine learning model has been developed, to this date, to approximate delivery costs for known districts with random distributions of requests. Moreover, no study has employed these techniques or measured the impact of different estimation approaches in the context of strategic problems such as districting-and-routing. The present paper fills these gaps.
@@ 3 Problem Statement
[…]
District costs. The long-term operational cost $\Phi(d)$ of any district $d$ is calculated as follows. Within each BU $i\in\{1,\dots,n\}$, we assume that random demand locations, represented as finite point sets $\mathbf{X}_{i}$, are generated according to the probability distribution $\mathbb{P}_{i}$ of a spatial Poisson process (see, e.g., Baddeley 2006) with intensity proportional to the population density. Consequently, demand requests are spread uniformly over the geographic area covered by each BU, and BUs with a larger population density have more requests on average per square kilometer. The cost of any district $d$ corresponds to the expected distance of the best tour leaving the depot, visiting the customer’s locations of all BUs $i\in d$, and returning to the depot. Therefore, the operational cost of $d$ is defined as:
$\Phi(d)=\mathbb{E}\left(\min_{\pi}\left\{\texttt{Dist}(\pi)|\pi\in\mathfrak{S}\left(\bigcup_{i\in d}\mathbf{X}_{i}\right)\right\}\right),$ (1)
where $\mathfrak{S}(X)$ is the set of all permutations of a set $X$, and $\texttt{Dist}(\pi)$ is the travel distance starting from the depot, visiting all locations according to the sequence $\pi$, and returning to the depot. With this definition, objective function $\Phi$ is monotonic, i.e., $\Phi(d^{\prime})\leq\Phi(d)$ if $d^{\prime}\subseteq d$. However, contrary to intuition, $\Phi$ is not submodular by direct consequence of the non-submodularity of optimal TSP distances (see, Anily and Federgruen 1990). Moreover, an exact evaluation of $\Phi$ is intractable as it requires calculating the expected optimal cost of a combinatorial optimization problem with stochastic parameters. To alleviate this issue, demand scenarios $X_{is}$ for $s\in S$ can be sampled from $\mathbb{P}_{i}$ for each BU $i\in d$, and the cost of the district can be approximated as:
$\Phi_{\textsc{saa}}(d)=\frac{1}{|S|}\sum_{s\in S}\min_{\pi}\left\{\texttt{Dist}(\pi)|\pi\in\mathfrak{S}\left(\bigcup_{i\in d}X_{is}\right)\right\}.$ (2)
@@
Graph partitioning formulation. The districting-and-routing problem can then be formally cast as a two-stage stochastic graph partitioning problem. Let $G(V,E)$ be an undirected graph, where each vertex $i\in V$ is a BU and edges $e\in E$ represent the contiguity between adjacent BUs (i.e., sharing a border). Let $\mathcal{D}=2^{V}$ be the set of all possible feasible districts respecting size and connectivity constraints. Then, the problem is formulated as the following integer program:
@@
For each possible district $d$, binary variable $\lambda_{d}$ takes value $1$ if and only if $d$ is selected in the solution. Parameter $b_{id}=1$ if BU $i$ appears in district $d$, and 00 otherwise. Objective (3) calculates the total cost of the selected districts. Constraints (4) ensure that each BU appears in exactly one district, and Constraint (5) fixes the number of districts. This problem can be seen as a two-stage stochastic program in which districts must be formed in the first stage, and one route needs to be selected in each district to visit the revealed customers in the second stage.
Formulation (36) is essentially of descriptive use, as there are two main barriers to its direct solution. First, it includes an exponential number of variables $d\in\mathcal{D}$. Moreover, the calculation of $\Phi(d)$ is very challenging, even when using an approximation as in Equation (2). For these reasons, this formulation is only practical to obtain baseline solutions for small problems. For larger cases, more efficient methods are needed. Designing scalable solution methods requires (i) efficient algorithms to estimate district costs (discussed in Section 4), and (ii) efficient search strategies for the graph-partitioning problem (discussed in Section 5).

### v2

# Ferraz, Ahmed, Cappart, and Vidal

@@
Ferraz, Ahmed, Cappart, and Vidal
[…]
Deep Learning for Data-Driven Districting-and-Routing
@@
[…]
Department of Mathematics and Industrial Engineering, Polytechnique Montréal, Canada.
@@
Districting-and-Routing, Strategic Optimization, Stochastic Optimization, Supervised Learning, Deep Learning, Graph Neural Networks
@@ 1 Introduction
Districting is the process of partitioning a service region, represented as a collection of basic geographical units, into larger clusters called districts. This practice is ubiquitous in large-scale transportation and last-mile delivery systems for mail delivery (Bruno et al. 2021), home care services (Benzarti, Sahin, and Dallery 2013), and maintenance services (García-Ayala et al. 2016). A delivery policy in fixed districts has multiple benefits: allowing the separation and the aggregation of the requests in advance before all information is available, reducing the complexity of the task thanks to the decomposition of the routing optimization process, stimulating the familiarity of drivers and thus their efficiency within their respective geographical regions (Zhong, Hall, and Dessouky 2007), and increasing the satisfaction of customers thanks to a higher familiarity with their drivers (Kovacs et al. 2014).
@@
Districting decisions are strategic and involve significant financial and societal stakes. These decisions typically hold for months or years, whereas operational routes occur on a daily or weekly basis and are subject to variation. Optimizing or even evaluating districting decisions is a very complex task. Demands are uncertain and volatile, and routing cost evaluations typically translate into large-scale vehicle routing problems, which are time-consuming to solve and highly sensitive to the spatial distribution of the requests. Because of these two different classes of decisions and planning horizons, the related districting-and-routing problems still represent important challenges (Drexl and Schneider 2015, Kalcsics and Ríos-Mercado 2019).
@@
An accurate method to estimate costs over a long planning horizon is to consider each district, use a Sample Average Approximation approach (SAA – Verweij et al. 2003) to generate sample demand scenarios (e.g., 10 to 500) within it, and solve a routing problem for each scenario and district. This approach offers the benefit of obtaining a good estimation, but at the cost of extensive computational time due to the large number of districts and scenarios. It is typically applicable to a fixed districting solution but becomes impractical within search algorithms for districting problems that require evaluating numerous districting solutions (e.g., a local search). Therefore, approaches for districting-and-routing regularly rely on continuous approximation formulas to evaluate costs (Franceschetti, Jabali, and Laporte 2017) without extensive computational overhead. It is, for instance, the case of the Beardwood-Halton-Hammersley (BHH) formula for the traveling salesman problem (Beardwood, Halton, and Hammersley 1959), which is frequently applied to estimate routing costs through $n$ independently uniformly distributed points in a compact area of size $A$ as $\alpha\sqrt{nA}$, where $\alpha$ is a constant. Though extremely fast, this evaluation approach is far less accurate than a scenario-based approach and may even drive the search toward solutions that have sub-optimal characteristics (see Section 6.4).
@@
Against this background, we leverage the recent progress in deep learning algorithms, specifically Graph Neural Networks (GNN – Scarselli et al. 2009, Kipf and Welling 2017), to accurately model expected routing costs and investigate their impact within solution methods for strategic districting-and-routing. Predictions with this approach should be fast enough to support solution-cost evaluations within a local search algorithm, and accurate enough to drive the search toward near-optimal districting solutions. Geographical areas are divided into geographical units called Basic Units (BU) and represented as a network of BUs with edges linking contiguous areas. Therefore, a solution to the districting problem corresponds to a partition of this graph into connected components representing contiguous districts. Our GNN learns to approximate the routing costs associated with any connected component assimilated to a district. It is trained on a set of districts with known costs obtained using SAA, i.e., by solving TSPs for demand scenarios with the Lin-Kernighan algorithm (Lin and Kernighan 1973, Helsgaun 2000) and taking their average value as the expected cost.
@@
The trained GNN gives us an oracle that estimates routing costs within districts accurately, as demonstrated by numerical experiments using out-of-sample districts. Furthermore, we must validate whether this performance translates into better strategic districting decisions, since even precise cost estimators may perform badly when integrated within an optimization process due to skew or outliers. To obtain a complete picture, we analyze the impact of using the proposed GNN within search algorithms for districting-and-routing. To that end, we design a prototypical solution approach for districting, which creates initial districts using a mathematical programming method for graph partitioning, and then subsequently applies an Iterated Local Search (ILS – Lourenço, Martin, and Stützle 2019) to generate better districting plans. The solution approach can use the GNN or any other baseline approach as a cost oracle, so we can conduct extensive computational experiments to compare the impact of different cost estimators on the structure and quality of the final solutions. Overall, the contributions of this paper are fourfold:
@@
    We design an ILS that leverages the predicted routing costs to generate high-quality districting plans. The construction of the initial solution is done by solving an adapted flow formulation for the balanced connected $k$-partition problem, whereas the local search and perturbation procedures are based on BU relocations and exchanges. Our method is generic and can use any cost oracle; it can produce efficient districting plans for practitioners and allows us to gauge the performance of accurate routing-cost estimates in the solution process.
@@
    We conduct extensive numerical experiments using the data of five metropolitan areas in the United Kingdom (Bristol, Manchester, Leeds, London, and the West Midlands), which contain up to 120 basic geographical units. We compare different cost oracles regarding the quality of routing cost estimates and their performance on the districting task, with districts of different sizes, instances of different scales, and different depot configurations. As seen in these experiments, the proposed GNN-based approach achieves a much smaller prediction error on average compared to other baselines, allowing long-term economic gains of $10.12\%$ on average (and exceeding $20\%$ in some cases) due to better districting decisions.
@@
    We investigate why cost estimation accuracy has a much larger impact than initially anticipated on the districting-and-routing problem. Indeed, whereas cost estimation accuracy differs by one or two percentage points between methods, their impact on districting-and-routing solution costs is one order of magnitude larger. A finer-grained analysis of the districting solutions shows that classical cost-estimation methods tend to drive the search toward suboptimal solution structures, and that the compactness of the districts is not sufficient to warrant good districting-and-routing solutions in all situations.
@@ 2 Related Works
Districting problems arise in numerous contexts in which a territory must be partitioned into smaller, connected regions. Comprehensive reviews are available in Zoltners and Sinha (2005) and Kalcsics and Ríos-Mercado (2019), which classify applications ranging from political redistricting to sales, health-care, and logistics planning. In political districting, compactness and fairness are often prioritized (see, e.g., Horn, Hampton, and Vandenberg 1993, Webster 2013). In contrast, operational districting, relevant for logistics, delivery, and health-care services, focuses primarily on workload balance, connectivity, and travel cost efficiency. This diversity of objectives has led to a broad spectrum of modeling and algorithmic approaches. To cite a few representative examples related specifically to logistics settings, Bozkaya, Erkut, and Laporte (2003) introduced a tabu search metaheuristic to solve a multi-criteria districting problem considering compactness and population equality, among other objectives. For home health care districting, Benzarti, Sahin, and Dallery (2013) introduced and solved different mixed integer linear programming (MILP) formulations, whereas García-Ayala et al. (2016) proposed an arc-based approach for problems with an underlying network structure, including postal delivery, meter readings, winter gritting, road maintenance, and municipal solid waste collection. For large-scale delivery systems, Bruno et al. (2021) analyzed the reorganization of postal collection zones in Bologna to adapt to declining mail volumes, combining spatial and operational constraints. Finally, in the sales and logistics domain, Lei et al. (2015) studied a multi-period districting-and-routing problem with multiple depots and proposed an adaptive large neighborhood search integrating equity, dissimilarity, and profitability objectives.
Long-term routing cost estimations in districting. In distribution contexts, long-term operational efficiency largely depends on the expected delivery costs within each district. Accurately evaluating these costs through a Sample Average Approximation (SAA) procedure (Verweij et al. 2003), that is, solving multiple delivery optimization problems such as Traveling Salesman Problems (TSP) or Vehicle Routing Problems (VRP) for sampled demand scenarios, provides reliable estimates but becomes prohibitively time-consuming when embedded in heuristic or metaheuristic frameworks exploring a large number of candidate districts.
To address this challenge, Continuous Approximation (CA) models (Franceschetti, Jabali, and Laporte 2017, Ansari et al. 2018) have long been employed as analytical surrogates for stochastic evaluations. They are widely used in districting and territory-design studies such as Daganzo (1984), Novaes, de Cursi, and Graciolli (2000), Galvão et al. (2006), Lei, Laporte, and Guo (2012) and Lei et al. (2015). The classical Beardwood-Halton-Hammersley (BHH) theorem (Beardwood, Halton, and Hammersley 1959) expresses the expected tour length through $R$ random points uniformly distributed over an area $A$ as $\alpha\sqrt{RA}$, with $\alpha$ an empirical constant depending on the distance metric. This relation has inspired numerous refinements. Daganzo (1984) extended it to vehicle-routing settings by adding a line-haul component proportional to the distance from the depot, while Chien (1992) performed Monte Carlo simulations to determine the best constant in different settings. Some subsequent studies sought to relax the uniformity assumptions: Figliozzi (2007) considered non-uniform demand densities and real urban network effects, and Çavdar and Sokol (2015) proposed a distribution-free regression model estimating TSP tour lengths using geometric dispersion and centrality measures. Finally, Franceschetti, Jabali, and Laporte (2017) extended CA formulations to heterogeneous, time-dependent, and network-based logistics systems. Although CA methods are extremely fast, they are fundamentally grounded on large-sample approximations and tend to lose accuracy on small or irregular instances. As will be shown later in this paper, their use within districting-and-routing optimization can also steer search heuristics toward geometrical characteristics that are not necessarily aligned with operational efficiency.
Data-driven estimators of routing costs. Data-driven research seeks to approximate routing distances from broader data and instance features. These approaches extend the principles of CA formulas, which can be interpreted as linear regression models built on a few spatial descriptors. Kwon, Golden, and Wasil (1995) presented an early use of shallow neural networks (with a single hidden layer of three neurons) to estimate the optimal TSP length for uniformly distributed customers in rectangular regions of up to 80 points. More recently, Akkerman and Mes (2025) developed linear models with 18 topological features to forecast routing costs in rolling-horizon logistics decisions, while Varol, Özener, and Albey (2024) proposed a three-layer neural network using engineered geometric descriptors of TSP instances and partial solutions generated by heuristics. Kou, Golden, and Poikonen (2022) and Kou, Golden, and Bertazzi (2024) introduced computational-statistics estimators based on the distribution of tour lengths obtained from one thousand randomized solutions. However, it is important to note that such estimators offer a practical advantage only when they are faster than solving the underlying problems directly. For medium-sized TSPs (around 100 customers), we observed that an efficient heuristic like Lin-Kernighan (LKH) algorithm (Helsgaun 2000) with a short termination criterion tends to be currently faster than the feature-extraction and inference procedures required by some of the aforementioned models.
Finally, progress has also been made on machine-learning approaches for combinatorial optimization that explicitly solve routing problems rather than estimate their cost. Reinforcement-learning and attention-based models have achieved good-quality solutions for medium-sized TSPs (see, e.g., Kool, van Hoof, and Welling 2018, Joshi, Laurent, and Bresson 2019, Kwon et al. 2020, Sun and Yang 2023, Drakulić, Michel, and Andreoli 2025). In contrast, other analyses highlighted the limited generalization of these approaches to larger graphs (Joshi et al. 2022). Graph Neural Networks (GNNs) play a central role in this research stream: various architectures (see, e.g., Dai, Dai, and Song 2016, Li et al. 2016, Hamilton, Ying, and Leskovec 2017) encode graph-structured data into latent embeddings that can be leveraged for decision making, and their use in combinatorial optimization has been comprehensively reviewed by Cappart et al. (2023).
Overall, despite major advances in deep learning, no advanced machine-learning model has been designed to approximate stochastic district delivery costs from BU-level characteristics, nor has any study examined how estimation strategies affect strategic districting-and-routing outcomes. Existing approaches span from fast but idealized CA formulas to learning-based estimators of deterministic routing costs and, more recently, neural architectures that directly solve routing problems (both of the latter needing repetitions on different demand scenarios to achieve reliable evaluation accuracy). Yet these methods are either not accurate or fast enough for large-scale districting optimization. This study addresses this gap by introducing a GNN-based methodology to estimate the expected routing cost of each district. Integrated within an ILS framework, this estimator permits rapid and accurate evaluation of numerous districting configurations, yielding high-quality districting-and-routing solutions. Furthermore, we provide a systematic analysis of how alternative cost-estimation models influence the structure and efficiency of strategic districting-and-routing decisions.
@@ 3 Problem Statement
We focus on a strategic districting problem encountered when organizing deliveries in a region from a central depot. Consider a geographical region consisting of $n$ geographical units called Basic Units (BU). We know the geographical boundaries and the population $\xi_{i}$ of each BU $i$. A district $d$ is a set of BUs, i.e., $d\,\subseteq\,2^{\Omega}$ with $\Omega=\{1,\dots,n\}$. The operational cost of a district is a function $\Phi:2^{\Omega}\to\mathbb{R}$ representing the expected (i.e., long-term) daily cost of delivering customers in the BUs of this district. Consequently, each district’s delivery operations are completely independent. The districting-and-routing problem studied in this paper then consists of partitioning the region into exactly $k$ districts in such a way that (i) each BU belongs to exactly one district, (ii) the number of BUs inside each district belongs to an admissible range $n_{\textsc{l}},n_{\textsc{u}}]$, (iii) the districts are connected in space, and (iv) the sum of the long-term operational costs of the districts is minimized.
District costs. The long-term operational cost $\Phi(d)$ of any district $d$ is calculated as follows. Within each BU $i$, we assume that a finite set $X_{i}(\omega)$ of random demand locations is generated according to the probability distribution $\mathbb{P}_{i}$ of a spatial Poisson process (see, e.g., Baddeley 2006) with intensity proportional to the population density. Consequently, demand requests are spread uniformly over the geographic area covered by each BU, and BUs with a larger population density have more requests on average per square kilometer. The cost of any district $d$ corresponds to the expected distance of the best tour leaving the depot, visiting the customer’s locations of all BUs $i\in d$, and returning to the depot. Therefore, the operational cost of $d$ is defined as:
$\Phi(d)=\mathbb{E}\left(\min_{\pi}\left\{\textrm{Dist}(\pi)|\pi\in\textrm{Perm}\left(\bigcup_{i\in d}X_{i}(\omega)\right)\right\}\right),$ (1)
where $\textrm{Perm}(X)$ is the set of all permutations of a set $X$, i.e., each permutation corresponds to a distinct sequence of the elements in $X$, and $\textrm{Dist}(\pi)$ is the travel distance starting from the depot, visiting all locations according to the sequence $\pi$, and returning to the depot. With this definition, objective function $\Phi$ is monotonic, i.e., $\Phi(d^{\prime})\leq\Phi(d)$ if $d^{\prime}\subseteq d$. However, an exact evaluation of $\Phi$ is intractable as it requires calculating the expected optimal cost of a combinatorial optimization problem with stochastic parameters. To alleviate this issue, we can sample $S$ demand scenarios $X_{is}$ for $s\in\{1,\dots,S\}$ from $\mathbb{P}_{i}$ for each BU $i\in d$, and approximate the cost of the district as:
$\Phi_{\textsc{saa}}(d)=\frac{1}{S}\sum_{s=1}^{S}\min_{\pi}\left\{\textrm{Dist}(\pi)|\pi\in\textrm{Perm}\left(\bigcup_{i\in d}X_{is}\right)\right\}.$ (2)
@@
Computational complexity. Due to the cardinality constraints on the number of BUs per district (between $n_{\textsc{l}}$ and $n_{\textsc{u}}$) and the requirement that the districts must remain connected, even identifying an initial feasible solution of the districting-and-routing problem is NP-hard by reduction from the balanced connected $k$-partition problem (Dyer and Frieze 1985). Moreover, computing the cost of the districts, even considering a single scenario, is another NP-hard subproblem.
Graph partitioning formulation. The districting-and-routing problem can be formally cast as a two-stage stochastic graph partitioning problem. Let $G(V,E)$ be an undirected graph, where each vertex $i\in V$ is a BU and edges $e\in E$ represent the contiguity between adjacent BUs (i.e., sharing a border). Let $\mathcal{D}\subset 2^{V}$ be the set of all possible feasible districts respecting size and connectivity constraints. Then, the problem is formulated as the following integer program:
@@
For each possible district $d$, binary variable $\lambda_{d}$ takes the value $1$ if and only if $d$ is selected in the solution. Parameter $b_{id}=1$ if BU $i$ appears in district $d$, and 0 otherwise. Objective (3) calculates the total cost of the selected districts. Constraints (4) ensure that each BU appears in exactly one district, and Constraint (5) ensures $k$ districts are formed. This problem can be seen as a two-stage stochastic program in which districts must be formed in the first stage, and one route needs to be selected in each district to visit the revealed customers in the second stage.
Formulation (36) is essentially of descriptive use, as there are two main barriers to its direct solution. First, it includes an exponential number of variables $d\in\mathcal{D}$. Moreover, the calculation of $\Phi(d)$ is very challenging, even when using an approximation as in Equation (2). For these reasons, this formulation is only practical for obtaining baseline solutions for small problems. For larger cases, more efficient methods are needed. Designing scalable solution methods requires (i) efficient algorithms to estimate district costs (discussed in Section 4), and (ii) efficient search strategies for the graph-partitioning problem (discussed in Section 5).